OpenStudy (anonymous):

Need serious help on this . . Simplify: tan40.45=x/150+(x/tan50) . . ANSWER=449.94m

OpenStudy (anonymous):

\[\tan40.45=\frac{ x }{ 150+\frac{ x }{ \tan50 } }\]

OpenStudy (badhi):

$$\begin{align*} \tan (40.45^\circ)&=\frac{x}{150+\frac{x}{\tan(50^\circ)}}\\ 150\tan(40.45^\circ)+\frac{\tan(40.45^\circ)}{\tan(50^\circ)}x&=x\\ 150\tan(40.45^\circ)&=\left(1-\frac{\tan 40.45^\circ}{\tan 50^\circ}\right)x\\ x&=\frac{150\tan 40.45^\circ}{\left(1-\frac{\tan 40.45^\circ}{\tan 50^\circ}\right)} \end{align*}$$

OpenStudy (anonymous):

so, @koli123able i got x = 1.00805